Find the slope of the line passing through the given points.
(-3, 3) and (3/4, -2)

Found 2 solutions by Alan3354, leon1068:
Answer by Alan3354(69443)

Given 2 points example.
Find the equation of the line thru the points (2,1) and (3,5)
---------------------
This is a 2 step process. First find the slope of the line thru the points.
slope, m = diffy/diffx
m = (5-1)/(3-2)
m = 4
---------
Now use y = mx + b with either point to find b, the y-intercept.
y = mx + b
5 = 4*3 + b
b = -7
-------
y = mx + b
y = 4x - 7 is the answer.
------------------------
For further assistance, or to check your work, email me via the thank you note.

Answer by leon1068(1)